OpenStudy (anonymous):

Shane pays $2568 per year in tuition. How much should he budget monthly for it? (Points : 3) $107 $156 $214 $228

OpenStudy (igreen):

12 months in a year. 2568 / 12 = ?

OpenStudy (love_ranaa):

107*12= 1284 156*12= 1872 214*12= 2568 228*12= 2736 And since he need 2568, he should budget how much :)